OpenStudy (cutiecomittee123):

which of the following are vertical asymptotes of y=2cot(3x)+4 x=0 x=2pi x=+-pi/2 x=pi/3

jimthompson5910 (jim_thompson5910):

cotangent is defined as cot(x) = cos(x)/sin(x)

jimthompson5910 (jim_thompson5910):

so, cot(3x) = cos(3x)/sin(3x)

jimthompson5910 (jim_thompson5910):

cot(3x) has vertical asymptotes when sin(3x) = 0 solve sin(3x) = 0 to get your answer

OpenStudy (cutiecomittee123):

how do I solve that?

OpenStudy (cutiecomittee123):

and I am confused on how that will help find the asymptotes

jimthompson5910 (jim_thompson5910):

well an alternative is to plug in the given choices to see whether they make sin(3x) equal to 0

jimthompson5910 (jim_thompson5910):

for example, choice A has x = 0 replace x with 0 to get... sin(3x) = sin(3*0) = sin(0) = 0 so when x = 0, the value of sin(3x) is equal to 0. All of this means x = 0 is a vertical asymptote of cot(3x)

jimthompson5910 (jim_thompson5910):

how about x = 2pi? what do you get when you plug that value in for x?